OpenStudy (anonymous):

Solve x^2 + 2x + 9 = 0

OpenStudy (anonymous):

by discrimante method

OpenStudy (bahrom7893):

D = b^2-4ac = 4 - 36 = -32

OpenStudy (anonymous):

Solving Quadratic Equations with Complex Numbers

OpenStudy (bahrom7893):

x1 = (-2 - 4*i*sqrt(2))/2

OpenStudy (bahrom7893):

x2 = (-2 + 4*i*sqrt(2))/2

OpenStudy (bahrom7893):

And when u pull out a two and cancel with the bottom: x1 = -1-2*i*sqrt(2) x2 = -1+2*i*sqrt(2)

OpenStudy (anonymous):

ohhhhhh ok

OpenStudy (bahrom7893):

the equation for x is: (-b - sqrt(D))/(2a)

OpenStudy (bahrom7893):

sqrt(D) = sqrt(-32) = sqrt(-1*16*2) = sqrt(-1)*sqrt(16)*sqrt(2) = i*4*sqrt(2) = 4i*sqrt(2)

OpenStudy (anonymous):

thank you